OpenStudy (anonymous):

Let x be a real number. Prove if x>1 then x^2>x.

OpenStudy (anonymous):

Induction might work here.

OpenStudy (oaktree):

Do you know what induction is, @zonazoo?

OpenStudy (anonymous):

yes, but I don't see how you would use that here.

OpenStudy (oaktree):

Okay, so let's try a different approach. Let's look at the behavior of x^2 for all x>1. Do you see anything useful you could use there?

OpenStudy (anonymous):

ummm, well, im not sure what you are looking for. if you square any number that is greater than 1, the answer will always be greater than the number you squared... the y value of the graph will always be greater on the quadratic than the linear function past 1. I would of thought you need to use contradiction, like suppose x<x^2. but again im really not sure what to do

OpenStudy (oaktree):

Contradiction would be a great way to prove this. Again, induction would be best, but contradiction would only be a line or two longer.

OpenStudy (anonymous):

im not sure how you would do either.

OpenStudy (oaktree):

Okay. So we have \[x^2 - x < 0\]So \[x(x-1) < 0\]But this means that exactly one of x or x-1 must be negative, since two negatives is positive and two positives is positive. Following so far?

OpenStudy (anonymous):

yes... and this method would be by contradiction correct.

OpenStudy (raden):

x > 1 multiply by x on both sides, we get x^2 > x :)

OpenStudy (oaktree):

Right. Because we're assuming that something is wrong at the beginning. Anyway, so can you think of what the next step might be?

OpenStudy (anonymous):

idk, just stating that either x<0 or x<1, which would be a contradiction since x>1???

OpenStudy (oaktree):

Kind of. Since exactly one of x, x-1 must be negative, we know that x-1 is -1 and x is 0. But that means that x is not greater than 1. Contradiction. QED. Get it?

OpenStudy (anonymous):

yes I do. What is QED?

OpenStudy (oaktree):

Quod erat demonstrandum. It's Latin for "what was required to be proved" - it's basically a formal way of saying that the proof is finished.

OpenStudy (anonymous):

oh okay. Thank you for all the help.

OpenStudy (oaktree):

No problem.

OpenStudy (anonymous):

I suppose induction may not be as easy as I had thought... the base case, namely, is giving me trouble. But if you restrict the domain here to be the natural numbers, it's easier to work out. Base case: \(x=2\). You have \(2^2=4>2\), so the inequality holds. Assume it holds for \(x=k\), so that \(k^2>k\). Next, show it holds for \(x=k+1\), i.e. \((k+1)^2>k+1\). The trick here is to write this inequality in terms of that given by the previous assumption: \[(k+1)^2=\color{red}{k^2}+\color{blue}{2k+1}>\color{red}{k}+\color{blue}{1}\] You know the red part on the left is greater than the red on the right, so now it's a matter of showing \(2k+1>1\), which is true for all \(k>0\). So the inequality holds for \(x=k+1\). But again, I'm not sure how one would adjust this for the real number system.